A particle moves in a straight line with constant acceleration under a constant force F? Relation between power and displacement and power and time

Grade:12

1 Answers

Vikas TU
7 years ago
in a straight line with constant acceleration under a constant force Fa = constant.
a = dv/dt
v = u + at
then,
Power = F x v = F(u+at).